OpenStudy (anonymous):

Thor invests some money in an account earning 8% interest, compounded continuously. After how many years will he double his investment? [Use the formla A = Pert and round answer to the nearest tenth.] A) 6.2 years B) 7.5 years C) 8.7 years D) 9.9 years

jimthompson5910 (jim_thompson5910):

A = P*e^(rt) 2P = P*e^(rt) ... replace A with 2P (investment P doubling) 2 = e^(r*t) 2 = e^(0.08*t) ... plug in the interest rate r = 8% = 0.08 use natural logs to isolate t